OpenStudy (juan1857):

Find an equation of the line that is tangent to the graph of f and parallel to the given line:

OpenStudy (juan1857):

\[Function: f(x)=\frac{ 1 }{ \sqrt{x} }\]

OpenStudy (juan1857):

\[Line:x+2y-6=0\]

OpenStudy (agent0smith):

First find the slope of the given line. Then find the derivative of f(x), and set it equal to that slope, solve for x. Then plug x back into f(x) and solve. Then you can make the equation of the line, using the slope from the first step, and the point (x, f(x))

OpenStudy (juan1857):

yeah Im having trouble setting it equal to the slope

OpenStudy (agent0smith):

Show work.

OpenStudy (juan1857):

\[x ^{-1/2}=\frac{ -x }{ 2 }\]

OpenStudy (juan1857):

that's the farthest I've gotten

OpenStudy (agent0smith):

That's... not close. You haven't even found the derivative of f(x). I've no idea what you did. Follow the steps in the order I gave. Show work.

OpenStudy (juan1857):